I've tried this problem and several like it many, many times. But I still cannot find a correct answer: One canned juice drink is 30% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 25L that is 6% orange juice?
Thank You

Answer by scott8148(5891)

let x=amount of 30% juice; so 25-x=amount of 5% juice

x(30%) + (25-x)(5%) = 25(6%)

multiplying by 100 to clear percentages ___ 30x + 125 - 5x = 150 ___ 25x = 25

x = 1

25-x = 24